OpenStudy (anonymous):

Find the limit of: Lim (X->9) (x^(3/2) - 3x) / (x^2 - 9x)

OpenStudy (anonymous):

\[Lim (x->9) \frac{ x ^{\frac{ 3 }{ 2 }} - 3x }{ x ^{2}-9x }\]

OpenStudy (anonymous):

did you consider that \(x^3=x\cdot x ^{\frac{1}{2}}\) ? this way you might to try to get rid of a \(x\) first by factoring it out. What will that do to your limit?

OpenStudy (anonymous):

excuse my type, I meant \(x^\frac{3}{2}=x \cdot x^\frac{1}{2}\)